OpenStudy (anonymous):

HOw to solve -100^(-3/2) ????????? Please help!

OpenStudy (anonymous):

\[-\frac{1}{1000} \]

OpenStudy (anonymous):

okay cool . But how did you get to there?

OpenStudy (anonymous):

Please show the steps :) thank yiou!!!

OpenStudy (anonymous):

You can break it into steps with parenthesis: \[-(((100^{1/2})^{3})^{-1}\] Evaluate starting at the innermost parenthesis: \[-(((10})^{3})^{-1}\] \[-(1000}^{-1}\] -(1/1000)

OpenStudy (anonymous):

why is it to the exponent 3 and the -1, and not just -3??

OpenStudy (anonymous):

\[-100{}^{\wedge}(-3/2)=-\frac{1}{\sqrt[2]{100^3}}=-\frac{1}{\sqrt[2]{1000000}}=-\frac{1}{1000} \]

OpenStudy (anonymous):

You can do the exponents in any order like (like robtobey).

OpenStudy (anonymous):

Thanks